# MoVA: Adapting Mixture of Vision Experts to Multimodal Context

### v1

@@ Abstract
As the key component in multimodal large language models (MLLMs), the ability of the visual encoder greatly affects MLLM’s understanding on diverse image content. Although some large-scale pretrained vision encoders such as vision encoders in CLIP and DINOv2 have brought promising performance, we found that there is still no single vision encoder that can dominate various image content understanding, e.g., the CLIP vision encoder leads to outstanding results on general image understanding but poor performance on document or chart content. To alleviate the bias of CLIP vision encoder, we first delve into the inherent behavior of different pre-trained vision encoders and then propose the MoVA, a powerful and novel MLLM, adaptively routing and fusing task-specific vision experts with a coarse-to-fine mechanism. In the coarse-grained stage, we design a context-aware expert routing strategy to dynamically select the most suitable vision experts according to the user instruction, input image, and expertise of vision experts. This benefits from the powerful model function understanding ability of the large language model (LLM) equipped with expert-routing low-rank adaptation (LoRA). In the fine-grained stage, we elaborately conduct the mixture-of-vision-expert adapter (MoV-Adapter) to extract and fuse task-specific knowledge from various experts. This coarse-to-fine paradigm effectively leverages representations from experts based on multimodal context and model expertise, further enhancing the generalization ability. We conduct extensive experiments to evaluate the effectiveness of the proposed approach. Without any bells and whistles, MoVA can achieve significant performance gains over current state-of-the-art methods in a wide range of challenging multimodal benchmarks. Codes and models will be available at https://github.com/TempleX98/MoVA.
Keywords:
Multimodal large language model Vision encoder Mixture-of-expert
00footnotetext: * Equal contribution. Corresponding authors.
@@ 1 Introduction
Significant achievements in multimodal large language models (MLLMs) have been witnessed due to their remarkable proficiency in solving open-world tasks. MLLMs acquire visual perception capacity while inheriting sophisticated reasoning abilities and knowledge from large language models (LLMs). The core idea behind MLLMs is projecting the vision encoder representation into an LLM through a projector, facilitating a general-purpose multimodal understanding.
@@
Table 1: Comparison of CLIP vs. state-of-the-art task-specific vision encoders. Our evaluation criteria encompass a variety of dimensions: comprehensive benchmarks [49], text-oriented Visual Question Answering (VQA) [56, 54], general VQA [23], object hallucination [39], Referring Expression Comprehension (REC) [81], Referring Expression Segmentation (RES) [81], and medical VQA benchmark SLAKE [42].
@@
We aim to start the exploration through empirical analysis of readily available vision experts. In particular, we focus on the multimodal capabilities of seven distinct state-of-the-art vision encoders based on LLaVA-1.5-7B [45]. The results in Tab. 1 reveal that MLLMs with these task-specific vision encoders achieve optimal performance in their respective area. Concurrently, we note that the plain fusion (concatenation) of vision encoder experts adopted in previous works (e.g., SPHINX) would not bring consistent improvement compared with the single task-specific vision expert in its proficient task. The inherent bias of each expert introduces biased information and leads to performance degradation in the plain fusion paradigm. For example, DINOv2 serves as an expert in visual grounding but performs poorly at text-oriented tasks. Representation of DINOv2 would be regarded as biased information in text-related scenarios so incorporating DINOv2 for these tasks would inevitably cause performance decrease. Consequently, a flexible method of vision encoder ensemble that dynamically activates and weights context-relevant task-specific vision experts can fully unleash the capacity of these models while avoiding model bias.
@@
In this paper, we propose MoVA, a powerful MLLM, adaptively routing and fusing task-specific vision experts with a coarse-to-fine mechanism. Inspired by the powerful tool-use capabilities of LLM[59], the coarse-grained context-aware expert routing aims to employ LLM to select vision experts with strong relevance to the user’s image and instruction from the expert model pool. To improve the efficiency and effectiveness of context-aware expert routing, we integrate expert-routing low-rank adaptation (LoRA) [21] into the LLM component of MoVA. The fine-grained expert fusion facilitates better extraction and integration of expert representations based on multimodal context. Specifically, the expert knowledge extractor in the mixture-of-vision-expert adapter (MoV-Adapter) will extract diverse task-specific knowledge from various vision experts through mixture-of-expert (MoE) cross-attention layers. The dynamic gating network can allocate precise expert-wise soft weights for the integration of extracted task-specific knowledge. Under the coarse-to-fine paradigm, we provide a flexible and effective manner of leveraging representation from experts based on multimodal context and model expertise, further enhancing the model generalization ability.
@@
We conduct comprehensive experiments on various benchmarks to evaluate the effectiveness of MoVA, including MLLM benchmarks, visual question answering (VQA), visual grounding, image segmentation, and biomedical understanding. Without any bells and whistles, MoVA can achieve significant performance gains over current state-of-the-art methods.
To sum up, the key contributions of this paper are as follows:
(1) By analyzing the performance of individual vision encoders versus the plain fusion of multiple encoders across various tasks, we reveal that the inherent bias of each vision encoder can diminish its generalization ability across other irrelevant domains.
@@
(2) We propose MoVA, a powerful MLLM composed of coarse-grained context-aware expert routing and fine-grained expert fusion with MoV-Adapter. Based on multimodal context and model expertise, MoVA fully leverages representation from multiple context-relevant vision encoder experts flexibly and effectively while avoiding biased information brought by irrelevant experts.
(3) We demonstrate the effectiveness of each component in MoVA by elaborate ablation studies. Without any bells and whistles, MoVA can achieve significant performance gains over current state-of-the-art methods in a wide range of challenging benchmarks.
@@ 2 Related Work
2.1 Large Language Models
Large language models (LLMs) have achieved remarkable progress in Natural Language Processing (NLP). The emergence of GPT-3 [7] demonstrates that models will manifest profound capabilities in few-shot learning and zero-short learning with increasing model parameters, training data, and training computation [27]. The powerful conversational and comprehension ability of ChatGPT and GPT4 [2] is also attributed to the generalization of LLMs. Meanwhile, many institutions are involved in the research on LLM pretraining and fine-tuning, bringing a series of open-source LLMs, including LLaMA [70, 71], Vicuna [13], Baichuan [78], Yi [1], Qwen [4], ChatGLM [14], InternLM [68], etc. Apart from the traditional dense causal transformer [72] paradigm, mixture-of-expert (MoE) is also a popular LLM architecture design. Switch Transformer [15] leverages sparse experts to scale up LLMs into trillion parameters, where a router will choose the most appropriate experts from the expert pool based on each input token. Considering only part of the experts will be involved in model training and inference, LLM with MoE can benefit from both large parameter complexity and low computation cost. Mistral $8$$\times$$7$B model outperforms the LLaMA2-70B model on multiple benchmarks, further verifying the effectiveness of the MoE.
2.2 Multimodal Large Language Models
Recent multimodal large language models (MLLMs) [37, 3, 47, 5, 12, 75, 73, 9, 63, 64] usually leverage the alignment from visual features to the linguistic feature space to achieve superior vision-language understanding capabilities based on off-the-shelf LLMs and vision encoders. CLIP vision encoder [60], which is trained in contrastive learning from billions of diverse image-text pairs [62, 61], is widely used among these works. For example, LLaVA [47] adopts an MLP projector to align visual tokens from the frozen CLIP vision encoder to the embedding layer of LLM. However, The representation from CLIP exhibits strong discriminative abilities in classification and recognition but only has limited performance on downstream tasks like location and relation understanding [22]. To break through this bottleneck, some works [12, 10] turn to unlock the CLIP vision encoder and further fine-tune the parameter with training data for downstream tasks. For instance, Qwen-VL [5] collected massive training data for grounding and OCR to jointly optimize the CLIP vision encoder and LLM. Recent works propose to involve an extra frozen vision encoder to enhance the performance of MLLMs. SPHINX [41] is one of the pioneers, where grounding capabilities have been significantly improved with the assistance of the DINOv2 [57]. Vary [75] introduces an extra encoder training on large-scale charts and document data to improve the performance on related downstream tasks.
@@ 3.1 Overview
MoVA comprises five key components: 1) a pre-trained large language model (LLM) that generates accurate responses given the image tokens and instructions; 2) a base vision encoder; 3) vision experts that generate task-specific vision latent features; 4) an auxiliary expert-routing low-rank adaption (LoRA) module that helps LLM select appropriate experts based on images and instructions; 5) mixture-of-vision-expert adapter (MoV-Adapter) that performs fine-grained expert fusion based on the multimodal context.
As illustrated in Fig. 1, MoVA consists of two stages: coarse-grained context-ware expert routing and fine-grained expert fusion with MoV-Adapter. First, our coarse-grained context-ware expert routing leverages the tool-use capabilities of LLM, routing the most appropriate experts from $N$ expert candidates via LLM to help the model answer the user’s question. We incorporate the expert-routing LoRA module into the LLM to improve the efficiency and effectiveness of expert routing. This expert-routing LoRA module is trained with expert routing annotations and can better align the LLM and the routing task. In the second stage, we turn to enhance the visual representation with a novel MoV-Adapter module in a fine-grained manner. More specifically, we leverage the cross-attention mechanism to extract the task-specific knowledge of representations from chosen experts. Meanwhile, the dynamic gating network in MoV-Adapter can allocate soft weights to the extracted knowledge of each expert according to the input image and instruction. Then the extracted knowledge can be effectively integrated into the foundational representation of the base vision encoder. Finally, the enhanced visual representation with instruction tokens is fed to the LLM to generate an accurate response. In Sec. 3.2 and Sec. 3.3, we will focus on our core contributions, the context-aware expert routing strategy, and the expert fusion with MoV-Adapter. In Sec. 3.4, we will introduce the training process.
@@
As a router, your task is to choose several models from a model pool to assist you. Below is a brief overview of the expertise of each model in the pool:
@@
Identify and select models that will best enable you to accurately answer questions. Please consider the image contents, questions, and expertise of these models when you perform selection. Answer with the model’s letter from the given choices directly.
@@
3.1.1 Pretrained Vision Encoders and LLM.
The vision encoders in MoVA consist of a base encoder and multiple task-specific vision encoder experts. We choose the pre-trained CLIP ViT-L-336px as the base encoder. Our vision experts include several state-of-the-art task-specific encoders: DINOv2, Co-DETR, SAM, Pix2Struct, Deplot, Vary, and BiomedCLIP. The corresponding expertise is presented in Tab. 1. For example, both Pix2Struct and Vary will be used when the user asks the MLLM to scan the document image. MoVA is flexible and easy to generalize to all decoder-only LLMs. We mainly consider Vicuna-7B and Yi-34B as our language model in this work.
3.2 Coarse-grained Context-aware Expert Routing
The context-aware expert routing strategy aims to employ the impressive reasoning capacity of LLM to select vision experts with strong relevance to the user’s image and instruction from the expert model pool.
3.2.1 Pipeline of Context-aware Routing.
Specifically, we perform the context-aware expert routing in three steps during inference. First, the input image, user questions, and descriptions of expert models are converted into appropriate instructions that prompt the MLLM to perform expert selection. An example of the prompt instruction input and selection output is shown in Tab. 2. Such a routing task does not require high-resolution input images, hence we directly downsample the base encoder’s visual feature to obtain a coarse image embedding (e.g., $64$ image tokens). Considering the difference between routing instructions and conventional multimodal data, avoiding task conflict is necessary. We integrate additional lightweight expert-routing LoRA layers into the LLM for accurate expert routing and further disentangle the routing task from original multimodal tasks (e.g., conversation about natural scenes). The downsampled image tokens and instruction tokens are then fed to the LLM as inputs. Finally, the LLM generates the output text and we parse it to determine which vision expert should be selected for fine-grained knowledge extraction in the second stage. For instance, as depicted in Tab. 2, the LLM directly outputs the option’s letter of DINOv2 and Pix2Struct, thus we only utilize them for the subsequent extraction. During training, we do not perform context-aware expert routing and replace the routing outputs with our routing annotations to improve efficiency.
3.2.2 Routing Data Construction.
@@ 3.2 Coarse-grained Context-aware Expert Routing
Compared with other MLLMs, MoVA requires additional routing annotations. We first introduce the formal definition of the data structure for an unambiguous understanding of the routing data. The data structure for expert routing introduces additional routing annotation $\mathcal{R}$ to the conventional multimodal data $(\mathcal{I},\mathcal{Q},\mathcal{A})$. Here, $\mathcal{I}$ represents the image, $\mathcal{Q}$ and $\mathcal{A}$ refer to the question-answer pair, and $\mathcal{R}$ refers to the expert set which contains the most appropriate ones to solve this question. Then the construction process for routing data can be formulated as $(\mathcal{I},\mathcal{Q},\mathcal{A})\rightarrow\mathcal{R}$, with the primary objective being to derive vision experts that optimally align with the sample $(\mathcal{I},\mathcal{Q},\mathcal{A})$. Intuitively, the language modeling loss can serve as an effective metric for evaluating how a data sample aligns with the vision expert. Specifically, we can reuse the LLaVA-1.5-7B models with various vision encoders presented in Sec. 1 to perform loss computation. Here, we denote the model with the base encoder as $\mathcal{M}_{0}$ and the model with $j$-th expert among $N$ experts as $\mathcal{M}_{j}$. For the $i$-th sample $(\mathcal{I}_{i},\mathcal{Q}_{i},\mathcal{A}_{i})$, we send it to models $\{\mathcal{M}_{j}|j\in\{0,1,\ldots,N\}\}$ and calculate the language modeling loss $\{\mathcal{L}_{i,j}|j\in\{0,1,\ldots,N\}\}$. The $j$-th expert is regarded as a useful expert for the $i$-th sample only if $\mathcal{L}_{i,j}<\mathcal{L}_{i,0}$ and will be added to the routing set $\mathcal{R}_{i}$. Note that we only keep up to 3 vision experts to avoid the increasing computation costs brought by too many additional experts. All the routing annotations of our training data are generated offline. We can directly parse and input these offline results to the subsequent expert fusion component and LLM during training.

### v2

@@ Abstract
As the key component in multimodal large language models (MLLMs), the ability of the visual encoder greatly affects MLLM’s understanding on diverse image content. Although some large-scale pretrained vision encoders such as vision encoders in CLIP and DINOv2 have brought promising performance, we found that there is still no single vision encoder that can dominate various image content understanding, e.g., the CLIP vision encoder leads to outstanding results on general image understanding but poor performance on document or chart content. To alleviate the bias of CLIP vision encoder, we first delve into the inherent behavior of different pre-trained vision encoders and then propose the MoVA, a powerful and novel MLLM, adaptively routing and fusing task-specific vision experts with a coarse-to-fine mechanism. In the coarse-grained stage, we design a context-aware expert routing strategy to dynamically select the most suitable vision experts according to the user instruction, input image, and expertise of vision experts. This benefits from the powerful model function understanding ability of the large language model (LLM). In the fine-grained stage, we elaborately conduct the mixture-of-vision-expert adapter (MoV-Adapter) to extract and fuse task-specific knowledge from various experts. This coarse-to-fine paradigm effectively leverages representations from experts based on multimodal context and model expertise, further enhancing the generalization ability. We conduct extensive experiments to evaluate the effectiveness of the proposed approach. Without any bells and whistles, MoVA can achieve significant performance gains over current state-of-the-art methods in a wide range of challenging multimodal benchmarks. Codes and models are available at https://github.com/TempleX98/MoVA.
@@ 1 Introduction
Significant achievements in multimodal large language models (MLLMs) [1, 2, 3, 4, 5, 6, 7] have been witnessed due to their remarkable proficiency in solving open-world tasks. MLLMs acquire visual perception capacity while inheriting sophisticated reasoning abilities and knowledge from large language models (LLMs) [8, 9, 10]. The core idea behind MLLMs is projecting the vision representation into an LLM through a projector, facilitating a general-purpose multimodal understanding.
@@
Table 1: Comparison of CLIP vs. state-of-the-art task-specific vision encoders. Our evaluation criteria encompass a variety of dimensions: comprehensive benchmarks [16], text-oriented Visual Question Answering (VQA) [17, 18], general VQA [19], object hallucination [20], Referring Expression Comprehension (REC) [21], Referring Expression Segmentation (RES) [21], and medical VQA benchmark SLAKE [22]. We use the same data for each model.
@@
We aim to start the exploration through empirical analysis of readily available vision experts. In particular, we focus on the multimodal capabilities of seven distinct state-of-the-art vision encoders based on LLaVA-1.5-7B [28]. The results in Table 1 reveal that MLLMs with these task-specific vision encoders achieve optimal performance in their respective area. Concurrently, we note that the plain fusion (concatenation) of vision encoder experts adopted in previous works [13] would not bring consistent improvement compared with the single task-specific vision expert in its proficient task. The inherent bias of each expert introduces biased information and leads to performance degradation in the plain fusion paradigm. For example, DINOv2 serves as an expert in visual grounding but performs poorly at text-oriented tasks. Representation of DINOv2 would be regarded as biased information in text-related scenarios so incorporating DINOv2 for these tasks would inevitably cause performance decrease. Consequently, a flexible method of vision encoder ensemble that dynamically activates and weights context-relevant task-specific vision experts can fully unleash the capacity of these models while avoiding model bias.
@@
In this paper, we propose MoVA, a powerful MLLM, adaptively routing and fusing task-specific vision experts with a coarse-to-fine mechanism. Inspired by the powerful tool-use capabilities of LLM [29], the coarse-grained context-aware expert routing aims to employ LLM to select vision experts with strong relevance to the user’s image and instruction from the expert model pool. Thanks to the strong generalization ability of LLM, we also can perform model routing for vision experts in open scenarios. The fine-grained expert fusion facilitates better extraction and integration of expert representations based on multimodal context. Specifically, the expert knowledge extractor in the mixture-of-vision-expert adapter (MoV-Adapter) will extract diverse task-specific knowledge from various vision experts through mixture-of-expert (MoE) cross-attention layers. The dynamic gating network can allocate precise expert-wise soft weights for the integration of extracted task-specific knowledge. Under the coarse-to-fine paradigm, we provide a flexible and effective manner of leveraging representation from experts based on multimodal context and model expertise, further enhancing the model generalization ability. As presented in Table 1, MoVA can preserve the optimal performance of a single relevant vision encoder by ignoring non-relevant experts on the GQA, POPE, and REC task. Besides, MoVA can further boost performances via the fine-grained fusion of multiple relevant vision experts on other tasks.
@@
We conduct comprehensive experiments on various benchmarks to evaluate the effectiveness of MoVA, including MLLM benchmarks, visual question answering (VQA), visual grounding, and biomedical understanding. Without any bells and whistles, MoVA can achieve significant performance gains over current state-of-the-art methods.
@@
The contributions of this work are three-fold: (i) By analyzing the performance of individual vision encoders versus the plain fusion of multiple encoders across various tasks, we reveal that the inherent bias of each vision encoder can diminish its generalization ability across other irrelevant domains. (ii) We propose MoVA, a powerful MLLM composed of coarse-grained context-aware expert routing and fine-grained expert fusion with MoV-Adapter. Based on multimodal context and model expertise, MoVA fully leverages representation from multiple context-relevant vision experts flexibly while avoiding biased information of irrelevant experts. (iii) We demonstrate the effectiveness of each component in MoVA by elaborate ablation studies. MoVA can achieve significant performance gains over state-of-the-art methods in a wide range of challenging benchmarks.
@@ 2 Related Work
Multimodal architectures [1, 3, 6, 30, 31, 32, 33, 34], optimization paradigm [35, 36], applications [37, 38, 39, 40], and benchmarks [41, 42, 43, 44, 45, 46, 47] have recently achieved remarkable progress and garnered unprecedented attention within the academic community. Multimodal large language models (MLLMs) usually leverage the alignment from visual features to the linguistic feature space to achieve superior vision-language understanding capabilities based on off-the-shelf LLMs and vision encoders. CLIP vision encoder [11], which is trained in contrastive learning from billions of diverse image-text pairs [48, 49], is widely used among these works. For example, LLaVA [3] adopts an MLP projector to align visual tokens from the frozen CLIP vision encoder to the embedding layer of LLM. However, The representation from CLIP exhibits strong discriminative abilities in classification and recognition but only has limited performance on downstream tasks like location and relation understanding [50]. To break through this bottleneck, some works [4, 51] turn to unlock the CLIP vision encoder and further fine-tune the parameter with training data for downstream tasks. For instance, Qwen-VL [6] collected massive training data for grounding and OCR to jointly optimize the CLIP vision encoder and LLM. Recent works propose to involve an extra frozen vision encoder to enhance the performance of MLLMs. SPHINX [13] is one of the pioneers, where grounding capabilities have been significantly improved with the assistance of the DINOv2 [15]. Vary [12] introduces an extra encoder training on large-scale charts and document data to improve the performance on related downstream tasks.
@@ 3.1 Overview
MoVA comprises five key components: (i) a pre-trained large language model (LLM) that generates accurate responses given the image tokens and instructions; (ii) a base vision encoder; (iii) vision experts that generate task-specific vision latent features; (iv) mixture-of-vision-expert adapter (MoV-Adapter) that performs fine-grained expert fusion based on the multimodal context.
As illustrated in Figure 1, MoVA consists of two stages: coarse-grained context-ware expert routing and fine-grained expert fusion with MoV-Adapter. First, our coarse-grained context-ware expert routing leverages the tool-use capabilities of LLM, routing the most appropriate experts from $N$ expert candidates via LLM to help the model answer the user’s question. In the second stage, we turn to enhance the visual representation with a novel MoV-Adapter module in a fine-grained manner. More specifically, we leverage the mixture-of-expert (MoE) cross-attention layers to extract the task-specific knowledge of representations from chosen experts. Meanwhile, the dynamic gating network in MoV-Adapter can allocate soft weights to the extracted knowledge of each expert according to the input image and instruction. Then the extracted knowledge can be effectively integrated into the foundational representation of the base vision encoder. Finally, the enhanced visual representation with instruction tokens is fed to the LLM to generate an accurate response. In Section 3.2 and Section 3.3, we will focus on our core contributions, the context-aware expert routing strategy, and the expert fusion with MoV-Adapter. In Section 3.4, we will introduce the training process.
Pretrained Vision Encoders and LLM. The vision encoders in MoVA consist of a base encoder and multiple task-specific vision encoder experts. We choose the pre-trained CLIP ViT-L-336px as the base encoder. Our vision experts include several state-of-the-art task-specific encoders: DINOv2, Co-DETR, SAM, Pix2Struct, Deplot, Vary, and BiomedCLIP. The corresponding expertise is presented in Table 1. For example, both Pix2Struct and Vary will be used when the user asks the MLLM to scan the document image. MoVA is flexible and easy to generalize to all decoder-only LLMs. We mainly consider Vicuna-7B [8], Llama3-8B  [^1] https://github.com/meta-llama/llama3, and Yi-34B [52] as our language models in this work.
3.2 Coarse-grained Context-aware Expert Routing
Pipeline of Context-aware Routing. The context-aware expert routing strategy aims to employ the impressive tool-use capacity of LLM to select vision experts with strong relevance to the user’s image and instruction from a model pool. Specifically, we perform the context-aware expert routing in three steps during inference. First, the input image, user questions, and descriptions of expert models are converted into appropriate instructions that prompt the MLLM to perform expert selection. An example of the prompt instruction input and selection output is shown in Table 2. Such a routing task does not require image details and high-resolution input images, hence we directly downsample the base encoder’s visual feature to obtain a coarse image embedding (e.g., $144$ image tokens). The downsampled image tokens and instruction tokens are then fed to the LLM as inputs. Finally, the LLM generates the output text and we parse it to determine which vision expert should be selected for fine-grained knowledge extraction in the second stage. For instance, as depicted in Table 2, the LLM directly outputs the option’s letter of DINOv2 and Pix2Struct, thus we only utilize them for the subsequent extraction. During training, we do not perform context-aware expert routing and replace the routing outputs with our routing annotations to improve efficiency.
@@ 3.2 Coarse-grained Context-aware Expert Routing
You are a helpful assistant router. Based on the visual content, questions, and model pool the user provides, you need to consider the expertise of these models to select the most 3 suitable models to help you answer the questions. Answer with the model’s letter from the given choices directly. If no models are selected, just answer ’none’.
Model pool:
@@
Routing Data Construction. Compared with other MLLMs, MoVA requires additional routing annotations. We first introduce the formal definition of the data structure for an unambiguous understanding of the routing data. The data structure for expert routing introduces additional routing annotation $\mathcal{R}$ to the conventional multimodal data $(\mathcal{I},\mathcal{Q},\mathcal{A})$. Here, $\mathcal{I}$ represents the image, $\mathcal{Q}$ and $\mathcal{A}$ refer to the question-answer pair, and $\mathcal{R}$ refers to the expert set which contains the most appropriate ones to solve this question. Then the construction process for routing data can be formulated as $(\mathcal{I},\mathcal{Q},\mathcal{A})\rightarrow\mathcal{R}$, with the primary objective being to derive vision experts that optimally align with the sample $(\mathcal{I},\mathcal{Q},\mathcal{A})$. Intuitively, the language modeling loss can serve as an effective metric for evaluating how a data sample aligns with the vision expert. Specifically, we can reuse the LLaVA-1.5-7B models with various vision encoders presented in Section 1 to perform loss computation. Here, we denote the model with the base encoder as $\mathcal{M}_{0}$ and the model with $j$-th expert among $N$ experts as $\mathcal{M}_{j}$. For the $i$-th sample $(\mathcal{I}_{i},\mathcal{Q}_{i},\mathcal{A}_{i})$, we send it to models $\{\mathcal{M}_{j}|j\in\{0,1,\ldots,N\}\}$ and calculate the language modeling loss $\{\mathcal{L}_{i,j}|j\in\{0,1,\ldots,N\}\}$. The $j$-th expert is regarded as a useful expert for the $i$-th sample only if $\mathcal{L}_{i,j}<\mathcal{L}_{i,0}$ and will be added to the routing set $\mathcal{R}_{i}$. Note that we only keep up to 3 vision experts to avoid computation costs brought by too many additional experts. All the routing annotations of our training data are generated offline. We can directly parse and input these offline results to the subsequent expert fusion component during training.
Routing Data Augmentation. To preserve the expert routing robustness and generalization ability in open scenarios, we only randomly select 2K samples for training, remove the model name in model description, and rewrite the model descriptions using ChatGPT [53] for each expert. We also shuffle the model pool and randomly truncate the model pool during training.